Bob is moving and all of his CDs are mixed up in a box. Twelve CDs are rock, eight are jazz, and five are classical. If he reaches in the box and selects them at random, find each possibility.
A). P ( 1 classical, 2 jazz)
B). P ( 1 jazz, 2 rock)
C). P ( 2 rock, 2 classical)

Answer by stanbon(75887)

Bob is moving and all of his CDs are mixed up in a box. Twelve CDs are rock, eight are jazz, and five are classical. If he reaches in the box and selects them at random, find each possibility.
----
Rule for solving the problem:
Prob of an event = (# of ways to succeed)/(# of possible outcomes)
----------------------------------
A). P ( 1 classical, 2 jazz) = [5C1*8C2]/[25C3] = (5*28)/2300 = 140/2300
------------------------------------
B). P ( 1 jazz, 2 rock) = [8C1*12C2]/[25C3] = (8*66)/2300 = 528/2300
------------------------------------
C). P ( 2 rock, 2 classical) = [12C2*5C2]/[25C3] = (66*10)/2300 = 660/2300
=================
Cheers,
Stan H.